Refor to the data table bolow. Compiete parts (a) through (d). Click on the icon to view the data table. a. Find the bost rogrossion oquaton with IQ soore as the responso variable Use predictor variablos of brain volume (VOL) andlor wo ght (W) Se ect the correct choioe and fil in the answer boxes to complote your choice. (Round to four decimal places as needed.) 1Q, Brain Volume, and Weight Data Table Brain Volume Body Weight 57.372 8.233 64.947 8.038 63.704 61、043 132.661 107.831 62.423 82.975 61.215 1.135 84.327 9.097 97.565 99.773 81.395 8016 79.966 72.384 275 1051 174 106 1345 115 113 1102 158 Brain Volume Body Weight
Answer #1

##### R-code##
iq=c(97,90,87,86,101,103,104,96,126,126,102,97,94,89,93,85,96,115,113,124)
iq
volume=c(1003,963,1037,1029,1282,1275,1051,1060,1034,1071,1174,1081,1069,1106,1345,1438,1031,1102,1205,1158)
volume
weight=c(57.372,58.233,64.947,58.036,63.764,61.043,132.661,107.831,62.423,62.975,61.215,61.135,84.327,79.097,97.565,99.773,81.395,88.016,79.966,72.384)
weight
fit=lm(iq~volume+weight)
summary(fit)

###Answer:
## the option B is correct.

> iq=c(97,90,87,86,101,103,104,96,126,126,102,97,94,89,93,85,96,115,113,124)
> iq
[1] 97 90 87 86 101 103 104 96 126 126 102 97 94 89 93 85 96 115 113
[20] 124
> volume=c(1003,963,1037,1029,1282,1275,1051,1060,1034,1071,1174,1081,1069,1106,1345,1438,1031,1102,1205,1158)
> volume
[1] 1003 963 1037 1029 1282 1275 1051 1060 1034 1071 1174 1081 1069 1106 1345
[16] 1438 1031 1102 1205 1158
> weight=c(57.372,58.233,64.947,58.036,63.764,61.043,132.661,107.831,62.423,62.975,61.215,61.135,84.327,79.097,97.565,99.773,81.395,88.016,79.966,72.384)
> weight
[1] 57.372 58.233 64.947 58.036 63.764 61.043 132.661 107.831 62.423
[10] 62.975 61.215 61.135 84.327 79.097 97.565 99.773 81.395 88.016
[19] 79.966 72.384
> fit=lm(iq~volume+weight)
> summary(fit)

Call:
lm(formula = iq ~ volume + weight)

Residuals:
Min 1Q Median 3Q Max
-16.803 -8.385 -4.601 7.188 23.719

Coefficients:
Estimate Std. Error t value Pr(>|t|)
(Intercept) 112.121794 29.242129 3.834 0.00133 **
volume -0.005953 0.025884 -0.230 0.82086
weight -0.055025 0.159775 -0.344 0.73478

the option B is correct.
Thus the regression equation is ,

IQ=112.121794+(-0.005953)VOL +(-0.0055025) W